I need to find the area of a square with all sides = sqrt of 3 plus sqrt of 5. What I mean is that the square has four equal sides measure sqrt of 3 + sqrt of 5. I am having a tough time with this one.
PS I just asked another question on radicals 5 minutes ago and I cannot open the solution to the problem.
I hope someone can help me. Thanks.

Answer by ptaylor(2198)

I need to find the area of a square with all sides = sqrt of 3 plus sqrt of 5. What I mean is that the square has four equal sides measure sqrt of 3 + sqrt of 5. I am having a tough time with this one.

The area of a square (A) equals S^2 where S is the length of one side:
A=(sqrt3+sqrt5)^2 and this equals (multiplying it out):
(sqrt3)(sqrt3)+(sqrt3)(sqrt5)+(sqrt3)(sqrt5)+(sqrt5)(sqrt5) and we have
NOTE 1: (sqrt3)(sqrt3)=3
NOTE 2: (sqrt3)(sqrt5)=sqrt(15)
NOTE 3: (sqrt5)(sqrt5)=5
Putting it all together, we have
3+2sqrt(15)+5 and this equals:
8+2sqrt(15)=
A=2(4+sqrt(15))
A=15.7459 sq units

ck by using a calculator

Hope this helps-----ptaylor